OpenStudy (graciek):

Which of these is a rational number? A.π B.sqrt3 C.sqrt2 D.1.3 Mrs. Clarkson designed a rectangular garden the length of 15 meters and width of 8 meter she plans to build a walkway through the garden from one corner to the next what measure is closest length of the diagonal walkway? A. 8 meters B. 11 meters C. 17 meters D. 23 meters Which number is a perfect cube? A. 3 B. 15 C. 64 D. 567 When adding sqrt 25 and -9, which type of number is the sum? A. Whole number B. Irrational number C. Integer D. Radical

OpenStudy (anonymous):

you have a lot of questions which one you want to do first?

OpenStudy (graciek):

the first one

OpenStudy (anonymous):

ok pick the terminating decimal, that is a "rational number"

OpenStudy (anonymous):

\(\pi\) is not a fraction, neither is the square root of a number that is not a perfect square like say \(4\) or \(25\)

OpenStudy (anonymous):

2. is just simple Pythagoras

OpenStudy (anonymous):

for the next one use pythagoras \[a^2+b^2=c^2\\ 8^2+15^2=c^2\] solve for \(c\)

OpenStudy (anonymous):

never heard of "simple pythagoras" before was that smart pythagoras's idiot brother?

OpenStudy (anonymous):

64 is a square number (8 × 8) and a cube number (4 × 4 × 4). This happens if and only if the number is a perfect sixth power.

OpenStudy (anonymous):

@satellite73 srry i didnt understood what u meant by "never heard of "simple pythagoras" before was that smart pythagoras's idiot brother?"

OpenStudy (alexandervonhumboldt2):

ROFL

candycove (candycove):

LOL

OpenStudy (anonymous):

as for the last one, what is \(\sqrt{25}\)?

OpenStudy (graciek):

5

OpenStudy (anonymous):

ok and \(5-9=?\)

OpenStudy (graciek):

-4

OpenStudy (anonymous):

right what kind of number is that?

candycove (candycove):

Satellite, how do you feel being the only person to have '100' SmartScore?

OpenStudy (anonymous):

sorta like simple simon

OpenStudy (anonymous):

@graciek do u know what type of number -4 is?

OpenStudy (anonymous):

if u try googling each term, it ll take u 2 min and u ll know what each term means :)

OpenStudy (graciek):

ok

OpenStudy (anonymous):

Counting numbers (also called whole number or natural numbers) are those used to count physical objects in the real world. They are integers that can be zero or positive. An Irrational Number is a real number that cannot be written as a simple fraction. An integer (from the Latin integer meaning "whole") is a number that can be written without a fractional component. For example, 21, 4, 0, and −2048 are integers. Radicals are used to express roots in Math.

OpenStudy (anonymous):

hope this helps:)